OpenStudy (anonymous):

Write the sum using summation notation, assuming the suggested pattern continues. -4 + 5 + 14 + 23 + ... + 131

OpenStudy (anonymous):

it is an A.P find the number of terms by the formula tn=a+(n-1) d tn=131 a=-4 d=5-(-4)=5+4=9 then sum \[Sn=\frac{ n }{ 2 }\left( a+tn \right)\]

OpenStudy (anonymous):

@surjithayer these are my answer choices

OpenStudy (anonymous):

find n first by the formula i gave in the beginning.

OpenStudy (anonymous):

131=-4+(n-1)9 131+4=9n-9 135+9=9n n=16

OpenStudy (anonymous):

\[\sum_{0}^{15}(-4+9n)\] 0,1,2,...,15 are 16 terms. when n=0,a=-4 when n=15,tn=-4+9*15=-4+135=131